# RAAD-LLM: Adaptive Anomaly Detection Using LLMs and RAG Integration

### v2

@@ Abstract
Anomaly detection in complex industrial environments poses unique challenges, particularly in contexts characterized by data sparsity and evolving operational conditions. Predictive maintenance (PdM) in such settings demands methodologies that are adaptive, transferable, and capable of integrating domain-specific knowledge. In this paper, we present RAAD-LLM, a novel framework for adaptive anomaly detection, leveraging large language models (LLMs) integrated with Retrieval-Augmented Generation (RAG). This approach addresses the aforementioned PdM challenges. By effectively utilizing domain-specific knowledge, RAAD-LLM enhances the detection of anomalies in time series data without requiring fine-tuning on specific datasets. The framework’s adaptability mechanism enables it to adjust its understanding of normal operating conditions dynamically, thus increasing detection accuracy. We validate this methodology through a real-world application for a plastics manufacturing plant and the Skoltech Anomaly Benchmark (SKAB). Results show significant improvements over our previous model with an accuracy increase from $70.7\%$ to $89.1\%$ on the real-world dataset. By allowing for the enriching of input series data with semantics, RAAD-LLM incorporates multimodal capabilities that facilitate more collaborative decision-making between the model and plant operators. Overall, our findings support RAAD-LLM’s ability to revolutionize anomaly detection methodologies in PdM, potentially leading to a paradigm shift in how anomaly detection is implemented across various industries.
@@ 2 Background
Figure 1: SPC technique of MAMR to set control limits for process stability in a query series $Q_{i}$. Figure A and Figure B are moving average and moving range, respectively. UCL is the defined upper control limit and LCL is the defined lower control limit. Series data points outside of control limits are deemed “out of statistical control” and are labeled as anomalous. Out of control points can be seen before line (1). Points between lines (1) and (2) represent a stable process. Points after line (2) also represent a stable process, however, they are trending towards out of control. These points, therefore, are potentially problematic. MEWMA is then applied to further refine these control limits.
@@ 2.1 Fundamental Concepts and Terminology
A process is said to be “in statistical control” if it is not experiencing out of control signals or significant variations beyond normal statistical variations [21]. Statistical process control (SPC) techniques are commonly used in manufacturing for monitoring sequential processes (e.g., production lines) to make sure that they work stably and satisfactorily [22]. In monitoring the stability of a process, SPC plays an essential role [23, 24]. The idea is that processes that are in statistical control are deemed to be stable processes [21]. For this work, stable processes form a baseline for normal process behavior. The selection of SPC techniques are use-case specific. For this work, moving average moving range (MAMR) is performed in phase I, after which multivariate exponentially weighted moving average (MEWMA) is applied as a phase II SPC implementation.
The univariate MAMR charts are plotted for each process variable in a time-series instance as shown in Figure 1. Upper (UCL) and lower (LCL) control limits for the moving average ($X$) and moving range ($mR$) charts are calculated as follows.
@@
MEWMA charts, a multivariate extension of the univariate exponentially weighted moving average (EWMA) charts [25], are plotted in the case that interrelationships exist among process variables and are used to further refine the control limits of MAMR. A MEWMA chart for a multivariate time-series instance is shown in Figure 2. The quantity plotted on the MEWMA chart is the MEWMA statistic which is calculated as follows.
$T_{i}^{2}=Z_{i}^{T}\Sigma^{-1}_{Z_{i}}Z_{i}$ (4)
where
$Z_{i}=\Lambda X_{i}+(1-\Lambda)Z_{i-1}$ (5)
where $Z_{i}$ is the $i^{th}$ EWMA vector, $X_{i}$ is the $i^{th}$ observation vector $i=1,2,...,n$, $Z_{0}=0$, $\Lambda=$ diag($\lambda_{1},\lambda_{2},...,\lambda_{p}$), $0<\lambda_{j}\leq 1$, $j=1,2,...,p$ and $\Sigma_{Z_{i}}$ is the covariate matrix of $Z_{i}$.
Smaller values of $\lambda$ are more effective in detecting small shifts in the mean vector, $Z_{i}$. The MEWMA chart gives an out-of-control signal as soon as $T_{i}^{2}>h_{4}$, where $h_{4}(>0)$ is chosen to achieve a specified in-control average run length [25]. However, inertia problems can delay reaction to a shift, so Hotelling’s $\chi^{2}$ limits should always be used to help prevent such delays [25]. Therefore, the UCL for the MEWMA statistics, is calculated based on a chi-squared distribution.
$UCL=n\cdot\chi^{2}_{1-\alpha,df=n}$ (6)
where $n$ is the number of process variable inputs.
Figure 2: MEWMA control chart for a multivariate time-series instance. MEWMA is applied after MAMR as a phase II SPC implementation to further refine control limits. UCL is the defined upper control limit. Points above the UCL are deemed “out of statistical control” and are additionally labeled as anomalous during phase II implementation. RAAD-LLM is applied to all points within the refined control limits to enhance anomaly detection.
@@ 4.1.3 Data Processing Workflow
Each $i^{th}$ series $Q_{i},i\in N$, is then processed using SPC techniques as shown in Figure 1. Time series points deemed “out of statistical control” are labeled as anomalous and filtered out of $Q_{i}$ before further processing. SPC is applied again after the first set of outliers (or anomalies) are removed. This is done to ensure extreme values do not affect control limits.
Unlike, AAD-LLM, MEWMA is applied as a phase II SPC implementation to further refine MAMR control limits. Therefore, it can be assumed that time series $Q_{i}$ represents a stable process. We use this assumption in initializing our comparison dataset $C_{i}$ as our baseline for normal behavior as explained in the next paragraph. The idea is that once the comparison dataset is initialized, the model then updates its understanding of normalcy as each new query window is ingested.
@@ 5.1 Data and Analysis for the PdM Use-Case
Figure 9: Processed sensor data from the SKAB dataset. The selected signals are preprocessed to include only those experiments that were 20 minutes in duration. Each signal begins in a non-anomalous experimental state and continues until the end of the experiment. Non-anomalous states are shown in blue and anomalous states are shown in orange. Processed signals are then input into the frameworks.
@@ 5.2 Data and Analysis for the SKAB Dataset
  • Pressure - Pressure in the loop after the water pump (Bar)
@@ 5.3 Evaluation of Model Performance
Model Accuracy Baseline=0.71 Precision Recall F1 score
@@
Model Accuracy Baseline=0.55 Precision Recall F1 score
@@
TABLE II: Average evaluation metrics over multiple model runs. The baseline model is one that predicts every observation to belong to the majority class. The RAG pipeline for both RAAD-LLM and RAAD-LLMv2 integrate a CSV-based knowledge base to dynamically retrieve relevant information for z-score comparisons, allowing for responses that reflect either exact matches for input values or the closest matches when exact values are not found. RAAD-LLMv2 integrates LlamaIndex for seamless retrieval of relevant domain knowledge from the vector store. Unlike RAAD-LLM, RAAD-LLMv2 adds only the retrieved content to the prompt before being fed forward through the frozen LLM.
@@
TABLE III: Best outlier detection scores for each anomaly detection method implemented on the SKAB dataset, sorted by F1 score [48]. A selection of NNs and ML based fault detection methods were chosen to compare on the benchmarks. RAAD-LLM and RAAD-LLMv2 metrics are averaged over multiple model runs. Multimodality allows for the enriching of input series data with semantics to enable more collaborative decision-making between the model and plant operators. For this work, multimodality refers to a model being optimized to detect anomalies across both time-series data and text. A model that requires no training or fine-tuning on the data it is applied to is considered transferable with zero-shot capabilities. Unlike all other methods, AAD-LLM, RAAD-LLM, and RAAD-LLMv2 are not trained or fine-tuned on the dataset they are applied to and are multimodal without requiring any additional strategies.
@@
The brief results are shown in Table II. With $95\%$ confidence, for the use-case dataset, RAAD-LLM achieved an accuracy of $89.1\pm 1.7\%$, which is a significant improvement over the baseline model. Furthermore, RAAD-LLM’s precision of $96.7\pm 0.04\%$, recall of $88.2\pm 2.3\%$ and F1 score of $92.2\pm 1.3\%$ are all notable improvements over the previous architecture. For the SKAB dataset, RAAD-LLM achieved an accuracy of $71.6\pm 0.4\%$, F1 score of $73.5\pm 0.8\%$, FAR of $42.1\pm 0.9\%$, and MAR of $11.4\pm 2.3\%$. As with the use-case dataset, all evaluation metrics for the SKAB dataset show a significant improvement over the previous architecture.

### v3

@@ Abstract
Anomaly detection in complex industrial environments poses unique challenges, particularly in contexts characterized by data sparsity and evolving operational conditions. Predictive maintenance (PdM) in such settings demands methodologies that are adaptive, transferable, and capable of integrating domain-specific knowledge. In this paper, we present RAAD-LLM, a novel framework for adaptive anomaly detection, leveraging large language models (LLMs) integrated with Retrieval-Augmented Generation (RAG). This approach addresses the aforementioned PdM challenges. By effectively utilizing domain-specific knowledge, RAAD-LLM enhances the detection of anomalies in time series data without requiring fine-tuning on specific datasets. The framework’s adaptability mechanism enables it to adjust its understanding of normal operating conditions dynamically, thus increasing detection accuracy. We validate this methodology through a real-world application for a plastics manufacturing plant and the Skoltech Anomaly Benchmark (SKAB). Results show significant improvements over our previous model with an accuracy increase from $70.7\%$ to $88.6\%$ on the real-world dataset. By allowing for the enriching of input series data with semantics, RAAD-LLM incorporates multimodal capabilities that facilitate more collaborative decision-making between the model and plant operators. Overall, our findings support RAAD-LLM’s ability to revolutionize anomaly detection methodologies in PdM, potentially leading to a paradigm shift in how anomaly detection is implemented across various industries.
@@ 2 Background
Figure 1: SPC technique of MAMR to set control limits for process stability in a query series $Q_{i}$. Figure A and Figure B are moving average and moving range, respectively. UCL is the defined upper control limit and LCL is the defined lower control limit. Series data points outside of control limits are deemed “out of statistical control” and are labeled as anomalous. Out of control points can be seen before line (1). Points between lines (1) and (2) represent a stable process. Points after line (2) also represent a stable process, however, they are trending towards out of control. These points, therefore, are potentially problematic. RAAD-LLM is applied to all points within control limits to enhance anomaly detection.
@@ 2.1 Fundamental Concepts and Terminology
A process is said to be “in statistical control” if it is not experiencing out of control signals or significant variations beyond normal statistical variations [21]. Statistical process control (SPC) techniques are commonly used in manufacturing for monitoring sequential processes (e.g., production lines) to make sure that they work stably and satisfactorily [22]. In monitoring the stability of a process, SPC plays an essential role [23, 24]. The idea is that processes that are in statistical control are deemed to be stable processes [21]. For this work, stable processes form a baseline for normal process behavior. The selection of SPC techniques are use-case specific. For this work, moving average moving range (MAMR) is implemented.
The univariate MAMR charts are plotted for each process variable in a time-series instance as shown in Figure 1. This aspect of plotting and analyzing the MAMR charts for all process variables in parallel will cause an increase in the Type I error rate. Upper (UCL) and lower (LCL) control limits for the moving average ($X$) and moving range ($mR$) charts are calculated as follows.
@@ 4.1.3 Data Processing Workflow
Each $i^{th}$ series $Q_{i},i\in N$, is then processed using SPC techniques. For this work, the univariate MAMR charts are plotted for each process variable as shown in Figure 1. This aspect of plotting and analyzing the MAMR charts for all process variables in parallel will cause an increase in the Type I error rate. Time series points deemed “out of statistical control” are labeled as anomalous and filtered out of $Q_{i}$ before further processing. SPC is applied again after the first set of outliers (or anomalies) are removed. This is done to ensure extreme values do not affect control limits. Therefore, it can be assumed that time series $Q_{i}$ represents a stable process. We use this assumption in initializing our comparison dataset $C_{i}$ as our baseline for normal behavior as explained in the next paragraph. The idea is that once the comparison dataset is initialized, the model then updates its understanding of normalcy as each new query window is ingested.
@@ 5.1 Data and Analysis for the PdM Use-Case
Figure 8: Processed sensor data from the SKAB dataset. The selected signals are preprocessed to include only those experiments that were 20 minutes in duration. The first 3 minutes were discarded as process start-up. Each signal begins in a non-anomalous experimental state and continues until the end of the experiment. Non-anomalous states are shown in blue and anomalous states are shown in orange. Processed signals are then input into the frameworks.
@@ 5.2 Data and Analysis for the SKAB Dataset
  • Pressure - The pressure in the loop after the water pump (Bar)
@@ 5.3 Evaluation of Model Performance
Model Accuracy Precision Recall F1 score
Baseline 0.74 0.74 1.00 0.85
@@
Model Accuracy Precision Recall F1 score
Baseline 0.45 0.45 1.00 0.62
@@
TABLE II: Average evaluation metrics over the best 5 model runs. The baseline model is one that predicts every observation to belong to the positive class. The RAG pipeline for both RAAD-LLM and RAAD-LLMv2 integrate a CSV-based knowledge base to dynamically retrieve relevant information for z-score comparisons, allowing for responses that reflect either exact matches for input values or the closest matches when exact values are not found. RAAD-LLMv2 integrates LlamaIndex for seamless retrieval of relevant domain knowledge from the vector store. Unlike RAAD-LLM, RAAD-LLMv2 adds only the retrieved content to the prompt before being fed forward through the frozen LLM.
@@
TABLE III: Best outlier detection scores for each anomaly detection method implemented on the SKAB dataset, sorted by F1 score [48]. A selection of NNs and ML based fault detection methods were chosen to compare on the benchmarks. RAAD-LLM and RAAD-LLMv2 metrics are averaged over the best 5 model runs. Multimodality allows for the enriching of input series data with semantics to enable more collaborative decision-making between the model and plant operators. For this work, multimodality refers to a model being optimized to detect anomalies across both time-series data and text. A model that requires no training or fine-tuning on the data it is applied to is considered transferable with zero-shot capabilities. Unlike all other methods, AAD-LLM, RAAD-LLM, and RAAD-LLMv2 are not trained or fine-tuned on the dataset they are applied to and are multimodal without requiring any additional strategies.
@@
The brief results are shown in Table II. With $95\%$ confidence, for the use-case dataset, RAAD-LLM achieved an accuracy of $88.6\pm 2.1\%$, which is a significant improvement over the baseline model. Furthermore, RAAD-LLM’s precision of $92.6\pm 0.1\%$, recall of $91.1\pm 3.3\%$ and F1 score of $91.9\pm 1.7\%$ are all notable improvements over the previous architecture. For the SKAB dataset, RAAD-LLM achieved an accuracy of $71.6\pm 0.4\%$, F1 score of $73.5\pm 0.8\%$, FAR of $42.1\pm 0.9\%$, and MAR of $11.4\pm 2.3\%$. As with the use-case dataset, all evaluation metrics for the SKAB dataset show a significant improvement over the previous architecture.